Job-Order Costing: Calculating Unit Product Costs EXERCISE Applying overhead cost computingUnit cost Lo2-2, LO2-3 ewhard Company assigns overhead cost to jobs on the basis of 125% of direct labor cost. The job cost sheet for Job 313 includes $10,000 in direct materials cost and $12,000 in direct labor cost. A total of 1,000 units were produced in Job 313. Required: What is the total manufacturing cost assigned to Job 313? What is the unit product cost for Job 313?
Answer #1
Concepts and reason

Job order cost sheet: Job order cost sheet records the details about direct materials cost, direct labor cost, and manufacturing overheads. It shows the total cost of the job work and completion details.

Variable overheads costs: These costs vary directly proportionally to the level of activity. If there is a change in the activity level by 10% then there would be a change in the variable cost by 10% in either direction. Variable cost is directly proportional to the volume of the production. Variable cost per unit remains constant.

Fixed overheads costs: These costs do not vary with change in the level of activity. Fixed cost remains the same for the period. If there is an increase in the activity level, then the fixed cost per unit decreases and if there is a decrease in the activity level then the fixed cost per unit increases. Fixed cost per unit changes with the change in the activity level.

Pre-determined overhead rate: It is the rate which is calculated based on the activity base. This rate is used to allocate manufacturing overheads.

Fundamentals

Total job cost is calculated by adding the all cost incurred for the completion of the job i.e. direct materials cost, direct labor cost, and overheads cost.

Per unit product cost for the job is obtained by dividing the total job cost by the total units. Total overhead cost is the sum of fixed overhead cost and variable overhead cost.

Total direct materials are $10,000, total direct labor is $12,000, and the estimated overhead cost is 125% of direct labor.

Determine the total manufacturing cost assigned to the job as follows: -

ManufacturingCost=DirectMaterials+DirectLabor+OverheadsCost=$10,000+$12,000+$12,000×125%=$37,000\begin{array}{c}\\Manufacturing\,\,{\mathop{\rm Cos}\nolimits} t = \,Direct\,Materials + Direct\,Labor\, + Overheads\,{\mathop{\rm Cos}\nolimits} t\\\\ = \,\$ 10,000 + \$ 12,000 + \$ 12,000\, \times 125\% \\\\ = \$ 37,000\\\end{array}

Therefore, the total manufacturing cost assigned to the job is $37,000.

The total manufacturing cost assigned to the job is $49,500 and the total units are 1,000.

Determine the unit product cost for the job as follows: -

UnitProductCost=TotalJobCost÷TotalUnits=$37,000÷1,000=$37\begin{array}{c}\\Unit\,\Pr oduct\,\,{\mathop{\rm Cos}\nolimits} t\, = Total\,Job\,\,{\mathop{\rm Cos}\nolimits} t \div \,Total\,Units\\\\ = \,\$ 37,000 \div 1,000\\\\ = \$ 37\\\end{array}

Therefore, the unit product cost for the job is $37.

Ans:

The total manufacturing cost assigned to the job is $37,000.

The unit product cost for the job is $37.